3.
Butane with an evaporation temperature of 0°C is used as liquefied gas in camping stoves. Butane is also a great power for summer hikers in Finland as a gas for a camping stove. One summer day the temperature was + 15°C and the air pressure 0.95 bar. Calculate how many liters of carbon dioxide was formed when a hiker boiled a 450-gram bottle of butane gas. Suppose the combustion was complete and that the hiker managed to cook the bottle empty (in fact, the same pressure remained inside the bottle as it was outside the bottle).
1
Expert's answer
2020-12-28T04:41:52-0500

2C4H10 + 13O2= 8CO2 + 10H2O

M (C4H10) = 58.12 g/mol

T = 15 + 273 = 288 K

n (C4H10) = 450/58.12 = 7.74 g/mol

n (CO2) = 8/2 n (C4H10) = 8/2 x 7.74 = 30.96 mol

pV=nRT

R = 8.31 L⋅bar⋅K−1⋅mol−1

T = 15 + 273 = 288 K

V (CO2) = nRT/p = (30.96 x 3.14 x 288) / 0.95 = 29 471 L